# The Philosophical Foundations of Growing AI Like A Child

### v1

@@ 1 Introduction
Recent advancements in artificial neural networks, particularly transformer-based large language models (LLMs), mark perhaps the most impressive progress in artificial intelligence (AI). Never had any AI models perform so well on tasks that are known to require high-level reasoning abilities, not limited to solving challenging mathematical problems (Ahn et al., 2024), writing codes for complex programs that runs smoothly (Zhuo et al., 2024), drafting travel and business plans (Xie et al., 2024), and importantly, doing so all via natural languages which humans can effortlessly understand. However, despite such astonishing achievements, scientists and users have noticed limitations of LLMs that would seem bizarre for humans: their performances on these tasks could fall drastically over slight tweaks of details in the task conditions (Yuan et al., 2023).
@@
Said limitations, which concern the robustness of problem-solving, have posed odds over the reliability of deploying LLMs to handle real-life tasks, at least in ideally low-supervised manners (Mitchell & Krakauer, 2023). Interpretations over the underlying causes of such limitations vary, with a prominent account citing the distinction between formal and functional linguistic competences — the ability to produce fluent languages versus the ability to understand and use them in the real world — and arguing that the robustness challenge is due to LLMs possessing the former without the latter (Mahowald et al., 2024). The extreme version of this account would be the one claiming that LLMs lack understanding altogether. In other words, they are “stochastic parrots” that can only solve tasks by abusing spurious correlations in the existing dataset, thereby failing to sensibly answer when the questions are dissimilar enough to what they have been trained on (Bender et al., 2021).
@@
At the same time, a surge of attention has been put into benchmarking LLMs: assessing them on large-scale repositories of tasks systematically-developed to target distinct reasoning abilities. A key motivation for such efforts is the notoriously mysterious nature of how LLMs work. It is not apparent, even to engineers who built them, how next-token predictions could enable problem-solving that requires advanced inference. By applying experimental paradigms with controlled designs, researchers are able to differentiate what LLMs can and cannot do under specific task conditions. One particularly staggering kind of result from the benchmarking efforts is the discovery that LLMs could fail miserably at tasks that are easy to humans, despite their high achievements on much harder ones (Dentella et al., 2024). Most interestingly, such a discrepancy could occur between tasks demanding abilities on seemingly the same domain. For example, GPT-4o, a state-of-the-art model open to the public, was shown giving wrong answers to a range of basic mathematical questions, some as simple as 1-digit counting, while crushing advanced problem sets that are challenging even to professional mathematicians (Williams & Huckle, 2024).
@@
    First, it suggests that the robustness challenge and the Moravec’s paradox are two sides of the same coin: they both arise from the discrepancy between the cognitive development of humans and machines, which can be summarized as scaling up vs. growing up [^1] This statement of the distinction is likely used first by Josh Tenenbaum in some versions of the talk titled “Scaling AI the Human Way”. Earliest version containing the statement is possibly given at IPAM (UCLA; Nov. 8, 2024)..
@@
A widely held belief in the current AI research community is that changes in LLMs’ reasoning performance can be directly attributed to changes in their scale, as tracked by the number of parameters they have in the neural networks and the size of the dataset they are trained on. This belief, hailed as the scaling law, has largely been supported by empirical results observed throughout LLMs’ progression throughout the past several years, and continues to be the primary strategy adopted by major companies for developing more advanced models (Kaplan et al., 2020). In this sense, scaling law has been taken as a theory regarding the nature of LLMs’ cognitive development, that is, increased computational power supporting a domain-general learning mechanism (Long, 2024). As Richard Sutton, in his seminal The Bitter Lesson, puts it:
@@
“One thing that should be learned from the bitter lesson is the great power of general purpose methods, of methods that continue to scale with increased computation even as the available computation becomes very great…we should stop trying to find simple ways to think about the contents of minds, such as simple ways to think about space, objects, multiple agents, or symmetries. All these are part of the arbitrary, intrinsically-complex, outside world. They are not what should be built in, as their complexity is endless; instead we should build in only the meta-methods that can find and capture this arbitrary complexity.” ((Sutton, 2019), p.2; emphasis added)
However, a principal limitation of scaling law as a theory of cognitive development is that it offers minimal explanation: it is not apparent how a neural network can master complex reasoning abilities just “emerge” through scaling up its number of parameters and the size of the dataset. This is not just for abilities yet to be achieved, but applies as well to those that have already emerged in bigger LLMs but not smaller ones. As Churchland and Sejnowski (1990) noted in the early days of connectionist research, an epistemological theory must not solely rely on a priori philosophy but should be grounded upon empirical data across levels of analysis (Churchland & Sejnowski, 1990). Claiming that all current barriers between LLMs and human intelligence can be resolved by scaling them up simply because it worked well in the past, without knowing how and why it worked well, would be an a priori argument with little scientific merit (Marcus, 2018; Long, 2024).
@@
On the other hand, understanding how human cognition develops as we grow up, and eventually reaches the level of (adult) intelligence that AI strives to achieve, has been the objective of developmental psychologists over the past several decades (Spelke, 2023; Tomasello, 2024). The application of specialized experimental paradigms for probing cognitive abilities have produced a large amount of psychological and behavioral data regarding what children can and cannot do across different age groups, which fertilizes empirically-grounded theories of human cognitive development.
@@
While many details in Piaget’s theory have been debated following more rigorous empirical investigations, modern literature generally support his idea that children first learn simple abilities to reason about the physical world and develop more complex, abstract abilities on top as they grow up [^2] Most of the debates have been regarding the specific ages of which children acquire different abilities. The difficulty of reaching consensus across empirical sources have largely led to the abandonment of the “stage” notion even among theorists who support neo-Piagetian approaches (Rochat, 2024). It is, however, important to note that the lack of clearcut age parameters do not go against the overall idea that children transition across a hierarchy of simple to complex abilities throughout development — the transition may just not be in a stage-to-stage fashion where each stage is qualitatively different from another in a definable way. (Barsalou, 2008; Samuelson & Smith, 2000; Barsalou, 2010; Pezzulo et al., 2013). This claim of a grounding nature of human cognitive development offers critical insights to the limitations in LLMs. Foremost, having learned simple abilities prior to complex abilities, humans are not subjected to the Moravec’s Paradox [^3] Although, to be fair, the Moravec’s Paradox is termed so on the basis that human intelligence is held as a standard for comparison..
Moreover, since there is a causal relationships postulated between the primitive, simple abilities and the late-coming complex abilities, the grounding view implies that the absence of simple abilities at an earlier time frame would likely affect the complex abilities later, hence a mechanistic link between the Moravec’s Paradox and the robustness challenge, both of which are observed in LLMs. It therefore may be the case that LLMs fail to generalize across conditions like humans when solving complex reasoning tasks because the skills they employ are not grounded upon simpler abilities on relevant domains.
@@
Note that this explanation hinges on the hypothesis that LLMs’ computational architecture share some aspects with that of humans, at least to the extent that the basic patterns of grounding apply. This may not be the case if human-level intelligence, specifically robust high-level reasoning, is multiple-realizable, which many argued to be likely (Bechtel & Mundale, 1999; McGrath & Russin, 2024). Especially, be reminded of scaling law, it is often postulated that LLMs’ cognitive development is based on some kind of domain-general learning mechanism, whereas research into human cognition and its grounding seems to be highly domain-specific.
@@
Nevertheless, human cognitive development might be more mechanistically similar to that of LLMs than it would appear. It is a long tradition to think that what drives the transitions from simple to complex abilities are highly domain-general. Piaget proposed that cognitive development is fundamentally the project of assimilation and accommodation, which is fitting familiar stimuli to existing cognitive structures and reconstructing them to interpret novel stimuli, respectively. These are essentially data-driven processes that do not select for particular patterns of information. In particular, Piaget suggested that the major transition differentiating the concrete operational stage from the preoperational stage is not the acquisition of specific abilities like conservation and perspective-taking per se, but the ability to perform systematic mental operations over structured mental representations, which in turn support said abilities by enabling functions such as reversibility and simulation (Piaget, 1965). Jerry Fodor, in his seminal work The Language of Thought, likewise interpreted that:
@@
“…a reasonable account of the stages of cognitive development could be elaborated by referring to increases in the expressive power of such systems. What I think one cannot have, however, is that concept learning provides the mechanisms for the stage- to-stage transitions. That is, if the child’s cognitive development is fundamentally the development of increasingly powerful representational/conceptual systems, then cognitive development cannot be the consequence of concept learning.” ((Fodor, 1975), p.89; emphasis added)
@@
From this perspective, it is therefore reasonable to think about LLMs’ limitations by mapping their differences with human cognition given a possible similarity between the common, foundational mechanism of their cognitive development. This is the view that both the robustness challenge and the Moravec’s Paradox are due to a missing link between simple abilities and complex abilities. Further, given that humans supposedly acquire complex abilities not by concept learning but rely on domain-general improvements, it is possible that this missing link is not due to problems with the way of increasing representational powers (i.e. scaling), but the failure to acquire simple abilities in the first place. In the next section, we begin to discuss theories and empirical results supporting this hypothesis, taken from both humans and LLMs.
@@ 3 Core Knowledge
A large body of works in cognitive science have demonstrated that humans possess a basic understanding of several key domains of the world at a very young age, henceforth core knowledge (Spelke, 2003; Spelke & Kinzler, 2007). This set of knowledge is generally understood as consisting of simple principles regarding objects, actions, number, space, and social relations, including how they relate to each other. Core knowledge is essentially children’s “developmental start-up software” that enables them to navigate and learn about the rich and dynamic nature of the environment in their early life (Lake et al., 2017).
Recent behavioural data have further supported a “child as scientist” proposal, showing that children appear to actively formulate intuitive hypothesis and validate them using these abilities to derive knowledge regarding previously unknown aspects of the world (Gopnik, 1996; Schulz et al., 2007; Gweon et al., 2010; Köksal-Tuncer & Sodian, 2018). These data, while supporting Fodor’s view that hypothesis-testing plays a foundational role in learning during the early days, highlights the domain-specific nature of the process, during which core knowledge of different aspects of the world is employed by the representational resources of the child.
@@
It is unclear how core knowledge comes about in the minds of individuals. Neuroimaging studies have suggested that functions within these key domains are likely organized modularly even in the children’s brain, exhibiting specialized neural networks supporting core systems of numerical operations (number), spatial navigation (space), theory-of-mind (social relations), and so on (Siegal & Varley, 2002; Newcombe & Sluzenski, 2004; Nieder & Dehaene, 2009). This evidence, paired with behavioural findings showing that these core-level abilities are shared by even the most remote, independently-evolved cultural groups, appears to favor the view that core knowledge is somehow innate. In contrast, others have speculated that it might just be early-developed knowledge instead of being “hardwired” in the brain at birth (Carey, 2011). These discussions constitute a major venue of the modern nativist vs. empiricist debate that we would not go into in detail here. In any case, core knowledge appears so early in childhood that it certainly cannot be the result of language-based learning. In turn, languages are likely a set of complex abilities that is developed upon these simple abilities. This marks a major discrepancy between the growing up of humans and the scaling up of LLMs, which acquire all reasoning abilities through linguistic data. However, it would again be a priori to directly view this as an explanation against LLMs’ possession of simple abilities, which undermines the multiple-realizability of cognition. The rapid development of Multimodal LLMs (MLLMs), which are LLMs that can receive and process multimodal (mostly visual as in current models) input, further offers potentials toward this direction by supporting sensory-based learning within language model architectures.
@@
Whether LLMs could or already possess simple abilities like human core knowledge is fundamentally an empirical question. Fortunately, benchmarking approaches inspired by experimental paradigms in human research have provided a promising pathway toward answering such questions, in which LLMs are essentially evaluated the same way as human participants in psychology experiments (Binz & Schulz, 2023; Shiffrin & Mitchell, 2023). Following such approaches, a large-scale benchmark developed by (Li et al., 2025) assessed twelve different cognitive abilities in MLLMs. The abilities tested range from simple, core-level abilities like object permanence and perceptual continuity to complex abilities like tool-using and intentionality understanding, spanning over all core knowledge domains. Tasks in said benchmark consist primarily of image or video-formatted adaptations of classic cognitive tasks used in the developmental psychology literature, such as visual cliff task (spatiality), three-mountain task (perspective-taking), and gear system (mechanical reasoning). Remarkably, they found that MLLMs generally perform much better in simple abilities than complex ones. As an example, state-of-the-art models like GPT-4o, while achieving near-human performance on abilities such as intentionality understanding—a higher-level theory-of-mind ability that typically emerges in humans no earlier than 6–7 years of age—fail catastrophically in simpler abilities like level-2 perspective-taking, which children are reported to master as early as 36 months of age (Moll & Meltzoff, 2011; Linsley et al., 2024; Gao et al., 2024).
Moreover, further in-depth analyses have shown that model performance on low-level cognitive abilities, such as perspective-taking, does not improve alongside model size (Gao et al., 2024; Li et al., 2025). In other words, scaling does not seem to allow MLLMs to become better at such abilities. While still an ongoing effort, said results have begun to provide critical empirical support for the hypothesis that scaling up LLMs may not allow them to grasp simple abilities available to humans even in early childhood.

### v2

@@ 1 Introduction
Recent advancements in artificial neural networks, particularly transformer-based large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; Bai et al., 2023; Jaech et al., 2024), mark perhaps the most impressive progress in artificial intelligence (AI). Never have any AI models performed so well on tasks known to require high-level reasoning abilities, including solving challenging mathematical problems (Ahn et al., 2024b), writing code for complex programs that run smoothly (Zhuo et al., 2024), drafting travel and business plans (Xie et al., 2024), and, importantly, doing so entirely through natural language that humans can effortlessly understand. With the emergence of multi-modal language models (MLLMs) (Li et al., 2023; Liu et al., 2024; Gemini, 2023), which align sensory input with language representations, these models show promise for developing intelligence grounded in the physical world and have already demonstrated unprecedented abilities in high-level perception and reasoning (Fu et al., 2023; Yang et al., 2025). However, despite such astonishing achievements, scientists and users have observed limitations in LLMs that would seem bizarre for humans: their performance on these tasks can drop drastically with slight tweaks to task conditions (Yuan et al., 2023; Varma et al., 2024).
@@
Said limitations, which concern the robustness of problem-solving, have posed odds over the reliability of deploying LLMs to handle real-life tasks, at least in ideally low-supervised manners (Mitchell & Krakauer, 2023). Interpretations over the underlying causes of such limitations vary, with a prominent account citing the distinction between formal and functional linguistic competences — the ability to produce fluent languages versus the ability to understand and use them in the real world — and arguing that the robustness challenge is due to LLMs possessing the former without the latter (Mahowald et al., 2024). The extreme version of this account would be the one claiming that LLMs lack understanding altogether (Deng et al., 2025). In other words, they are “stochastic parrots” that can only solve tasks by abusing spurious correlations in the existing dataset, thereby failing to sensibly answer when the questions are dissimilar enough to what they have been trained on (Bender et al., 2021).
@@
At the same time, a surge of attention has been put into benchmarking LLMs: assessing them on large-scale repositories of tasks systematically-developed to target distinct reasoning abilities. A key motivation for such efforts is the notoriously mysterious nature of how LLMs work. It is not apparent, even to engineers who built them, how next-token predictions could enable problem-solving that requires advanced inference. By applying experimental paradigms with controlled designs, researchers are able to differentiate what LLMs can and cannot do under specific task conditions. One particularly staggering kind of result from the benchmarking efforts is the discovery that LLMs could fail miserably at tasks that are easy to humans, despite their high achievements on much harder ones (Dentella et al., 2024; Li et al., 2025a). Most interestingly, such a discrepancy could occur between tasks demanding abilities on seemingly the same domain. For example, GPT-4o, a state-of-the-art model open to the public, was shown giving wrong answers to a range of basic mathematical questions, some as simple as 1-digit counting, while crushing advanced problem sets that are challenging even to professional mathematicians (Williams & Huckle, 2024).
@@
    First, it suggests that the robustness challenge and the Moravec’s paradox are two sides of the same coin: they both arise from the discrepancy between the cognitive development of humans and machines, which can be summarized as scaling up vs. growing up [^1] See also Tenenbaum (2024)..
@@
A widely held belief in the current AI research community is that changes in LLMs’ reasoning performance can be directly attributed to changes in their scale, as tracked by the number of parameters they have in the neural networks and the size of the dataset they are trained on. This belief, hailed as the scaling law, has largely been supported by empirical results observed throughout LLMs’ progression throughout the past several years, and continues to be the primary strategy adopted by major companies for developing more advanced models (Kaplan et al., 2020). In this sense, scaling law has been taken as a theory regarding the nature of LLMs’ cognitive development, that is, increased computational power supporting a domain-general learning mechanism (Long, 2024). As Sutton (2019), in his seminal ”The Bitter Lesson”, puts it:
@@
“One thing that should be learned from the bitter lesson is the great power of general purpose methods, of methods that continue to scale with increased computation even as the available computation becomes very great…we should stop trying to find simple ways to think about the contents of minds, such as simple ways to think about space, objects, multiple agents, or symmetries. All these are part of the arbitrary, intrinsically-complex, outside world. They are not what should be built in, as their complexity is endless; instead we should build in only the meta-methods that can find and capture this arbitrary complexity.” (p.2; emphasis added)
However, a principal limitation of the scaling law as a theory of cognitive development is that it offers minimal explanatory power: it remains unclear how complex reasoning abilities could simply emerge by increasing a model’s parameters or dataset size. This concern applies not only to hypothetical future capabilities, but also to those already seen in larger models but absent in smaller ones. As Churchland & Sejnowski (1990) emphasized in early connectionist research, epistemological theories must be grounded in empirical data across multiple levels of analysis, not just abstract philosophical reasoning. As such, claiming that further scaling alone will bridge all remaining gaps between LLMs and human cognition without understanding how and why scaling worked in the past amounts to an uninformative a priori assumption with limited scientific merit (Marcus, 2018; Long, 2024). While interest in mechanistic interpretability is growing, the scaling law remains largely unexplained at the algorithmic level (Marr, 1982), raising concerns about its sufficiency as a scientifically-grounded pathway for AI development.
@@
On the other hand, understanding how human cognition develops as we grow up, and eventually reaches the level of adults, where AI also strives to achieve, has been the objective of developmental psychologists over the past several decades (Spelke, 2023; Tomasello, 2024). The application of specialized experimental paradigms for probing cognitive abilities have produced a large amount of psychological and behavioral data regarding what children can and cannot do across different age groups, which fertilizes empirically-grounded theories of human cognitive development.
@@
While many details in Piaget’s theory have been debated following more rigorous empirical investigations, modern literature generally support his idea that children first learn simple abilities to reason about the physical world and develop more complex, abstract abilities on top as they grow up [^2] Most of the debates have been regarding the specific ages of which children acquire different abilities. The difficulty of reaching consensus across empirical sources have largely led to the abandonment of the “stage” notion even among theorists who support neo-Piagetian approaches (Rochat, 2024). It is, however, important to note that the lack of clearcut age parameters do not go against the overall idea that children transition across a hierarchy of simple to complex abilities throughout development — the transition may just not be in a stage-to-stage fashion where each stage is qualitatively different from another in a definable way. (Barsalou, 2008; Samuelson & Smith, 2000; Barsalou, 2010; Pezzulo et al., 2013). This claim of a grounding nature of human cognitive development offers critical insights to the limitations in LLMs.
@@
Foremost, having learned simple abilities prior to complex abilities, humans are not subjected to the Moravec’s Paradox [^3] Although, strictly speaking, the Moravec’s Paradox is termed so on the basis that human intelligence is held as a standard for comparison.. Moreover, since there are causal relationship postulated between the primitive, simple abilities and the late-coming complex abilities, the grounding view implies that the absence of simple abilities at an earlier time frame would likely affect the complex abilities later, hence a mechanistic link between the Moravec’s Paradox and the robustness challenge, both of which are observed in LLMs. It therefore may be the case that LLMs fail to generalize across conditions like humans when solving complex reasoning tasks because the skills they employ are not grounded upon simpler abilities on relevant domains. Note that this explanation hinges on the hypothesis that LLMs’ computational architecture shares some aspects with that of humans, at least to the extent that the basic patterns of grounding apply. This may not be the case if human-level intelligence, specifically robust high-level reasoning, is multiple-realizable, which many argued to be likely (Bechtel & Mundale, 1999; McGrath & Russin, 2024; Li et al., 2025a). Especially, be reminded of scaling law, it is often postulated that LLMs’ cognitive development is based on some kind of domain-general learning mechanism, whereas research into human cognition and its grounding seems to be highly domain-specific.
@@
Nevertheless, human cognitive development might be more mechanistically similar to that of LLMs than it would appear. It is a long tradition to think that what drives the transitions from simple to complex abilities are highly domain-general. Piaget proposed that cognitive development is fundamentally the project of assimilation and accommodation, which is fitting familiar stimuli to existing cognitive structures and reconstructing them to interpret novel stimuli, respectively. These are essentially data-driven processes that do not select for particular patterns of information. In particular, Piaget suggested that the major transition differentiating the concrete operational stage from the preoperational stage is not the acquisition of specific abilities like conservation and perspective-taking per se, but the ability to perform systematic mental operations over structured mental representations, which in turn support said abilities by enabling functions such as reversibility and simulation (Piaget, 1965). In his seminal work ”The Language of Thought”, Fodor (1975) likewise interpreted that:
@@
“…a reasonable account of the stages of cognitive development could be elaborated by referring to increases in the expressive power of such systems. What I think one cannot have, however, is that concept learning provides the mechanisms for the stage-to-stage transitions. That is, if the child’s cognitive development is fundamentally the development of increasingly powerful representational/conceptual systems, then cognitive development cannot be the consequence of concept learning.” (p.89; emphasis added)
@@
From this perspective, it is therefore reasonable to think about LLMs’ limitations by mapping their differences with human cognition given a possible similarity between the common, foundational mechanism of their development. This is the view that both the robustness challenge and the Moravec’s Paradox are due to a missing link between simple abilities and complex abilities. Further, given that humans supposedly acquire complex abilities not by concept learning but rely on domain-general improvements, it is possible that this missing link is not due to problems with the way of increasing representational powers (i.e. scaling), but the failure to acquire simple abilities in the first place. In the next section, we begin to discuss theories and empirical results supporting this hypothesis, taken from both humans and LLMs.
@@ 3 Core Knowledge
A large body of works in cognitive science have demonstrated that humans possess a basic understanding of several key domains of the world at a very young age, henceforth core knowledge (Spelke, 2003; Spelke & Kinzler, 2007). This set of knowledge is generally understood as consisting of simple principles regarding objects, actions, number, space, and social relations, including how they relate to each other. Core knowledge is essentially children’s “developmental start-up software” that enables them to navigate and learn about the rich and dynamic nature of the environment in their early life (Lake et al., 2017). Recent behavioural data have further supported a “child as scientist” proposal, showing that children appear to actively formulate intuitive hypothesis and validate them using these abilities to derive knowledge regarding previously unknown aspects of the world (Gopnik, 1996; Schulz et al., 2007; Gweon et al., 2010; Köksal-Tuncer & Sodian, 2018). These data, while supporting Fodor’s view that hypothesis-testing plays a foundational role in learning during the early days, highlights the domain-specific nature of the process, during which core knowledge of different aspects of the world is employed by the representational resources of the child.
@@
It is unclear how core knowledge comes about in the minds of individuals. Neuroimaging studies have suggested that functions within these key domains are likely organized modularly even in a child’s brain, exhibiting specialized neural networks supporting core systems of numerical operations (number), spatial navigation (space), theory-of-mind (social relations), and so on (Siegal & Varley, 2002; Newcombe & Sluzenski, 2004; Nieder & Dehaene, 2009). This evidence, paired with behavioural findings showing that these core-level abilities are shared by even the most remote, independently-evolved cultural groups, appears to favor the view that core knowledge is somehow innate. In contrast, others have speculated that it might just be early-developed knowledge instead of being “hardwired” in the brain at birth (Carey, 2011). These discussions constitute a major venue of the modern nativist vs. empiricist debate that we would not go into in detail here. In any case, core knowledge appears so early in childhood that it certainly cannot be the result of language-based learning. In turn, languages are likely a set of complex abilities that is developed upon these simple abilities. This marks a major discrepancy between the growing up of humans and the scaling up of LLMs, which acquire all reasoning abilities through linguistic data. However, it would again be a priori to directly view this as an explanation against LLMs’ possession of simple abilities, which undermines the multiple-realizability of cognition (Cao, 2022). The rapid development of MLLMs further offers potentials toward this direction by supporting sensory-based learning within language model architectures.
@@
Whether LLMs could or already possess simple abilities like human core knowledge is fundamentally an empirical question. Benchmarking approaches inspired by experimental paradigms in human research have provided a promising pathway toward answering such questions, in which LLMs are essentially evaluated the same way as human participants in psychology experiments (Binz & Schulz, 2023; Shiffrin & Mitchell, 2023). Following such approaches, a large-scale benchmark developed by Li et al. (2025a) assessed twelve different cognitive abilities in MLLMs. The abilities tested range from simple, core-level abilities like object permanence and perceptual continuity to complex abilities like tool-using and intentionality understanding, spanning over all core knowledge domains. Tasks in said benchmark consist primarily of image or video-formatted adaptations of classic cognitive tasks used in the developmental psychology literature, such as visual cliff task (spatiality), three-mountain task (visual perspective-taking), and fluid mechanics (mechanical reasoning). Remarkably, they found that MLLMs generally perform much better in complex abilities than simple ones. As an example, state-of-the-art models like GPT-4o, while achieving near-human performance on abilities such as intentionality understanding, a higher-level theory-of-mind ability that typically emerges in humans no earlier than 6–7 years of age, fail catastrophically in simpler abilities like visual perspective-taking, which children are reported to master as early as 36 months of age (Moll & Meltzoff, 2011; Linsley et al., 2024; Gao et al., 2024; Wang et al., 2025). Moreover, further in-depth analyses have shown that model performance on low-level cognitive abilities, such as perspective-taking, does not improve alongside model size (Li et al., 2025a). In other words, scaling does not seem to allow MLLMs to become better at such abilities. Said results provide critical empirical support for the hypothesis that scaling up LLMs may not allow them to grasp core knowledge.